40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 11 18:20:59 2018, gautam, Update, Optical Levers, BS/PRM HeNe is dead BSOL_dead.png
    Reply  Sat Dec 15 20:04:03 2018, gautam, Update, IOO, TT1/TT2 stepping 
       Reply  Wed Dec 19 15:15:56 2018, gautam, Update, IOO, TT1/TT2 stepping 
Message ID: 14368     Entry time: Wed Dec 19 15:15:56 2018     In reply to: 14362
Author: gautam 
Type: Update 
Category: IOO 
Subject: TT1/TT2 stepping 

I removed the ND filter from the ETMYT camera to facilitate searching for a TRY beam. This should be replaced before we go back to high power.

ELOG V3.1.3-